OpenStudy (anonymous):

Given a and b are constants, and solving with respect to t, how would you solve this integral?

OpenStudy (anonymous):

\[\int\limits_{}^{}e^{(a+b)t}dt\]

OpenStudy (anonymous):

Would you be able to use u-substitution, making (a+b)t = u?

OpenStudy (anonymous):

u = (a+b)t du = a+bdt \[\frac{ 1 }{ a+b }du = dt\]

OpenStudy (anonymous):

\[\frac{ 1 }{ a+b } \int\limits_{}^{}e ^{u}du\]

OpenStudy (anonymous):

I'm not sure if this is right at all :|

OpenStudy (anonymous):

I'm just as confused as you are unfortunately.

OpenStudy (anonymous):

It's okay, thanks for trying though! (: